Using the method of substitution to solve the system of linear equations, what is the ordered pair for x-7y=6 4x-28y=-15
Answer by IFY(62)


x-7y=6 4x-28y=-15
X = 6 + 7Y
4 (6 +7Y )- 28Y = -15
24 +28Y - 28Y = -15
Y = -15 +24
Y = 9
X = 6 + 7* 9
X = 69

HOPE THIS HELPS